OpenStudy (anonymous):

] Solve the system of linear equations by graphing. Write the solution as an ordered pair (x, y). 2x - 3y = 2 y = 2x + 2Steps for solving this system:1. Write the first equation in slope-intercept form y = mx+b2. Graph both linear equations in the graphing calculator for y1 and y23. Find the intersection of the lines using the Intersection feature (2nd, Trace, 5) 4. IF they intersect, then write the solution as an ordered pair (x, y)5. If the lines do not intersect and are parallel, then there is no solution. If they are the same line, then there are infinite solutions. *** intersect in 1

OpenStudy (anonymous):

Which steps of those you posted do you not understand? We can help you with the steps you are confused about.

OpenStudy (anonymous):

You are given 2x - 3y = 2 y = 2x + 2S First step is asking you to solve for y. Or in other words get y by itself.

OpenStudy (anonymous):

what is 2x-3y=2 in y=mx+b

OpenStudy (anonymous):

Sorry solve get this equation with y by itself 2x - 3y = 2

OpenStudy (anonymous):

All it's asking you to do is get y by itself. Remember we did this once,

OpenStudy (anonymous):

y=-2/3+2x?

OpenStudy (anonymous):

2x - 3y = 2 2x-2x - 3y = 2-2x - 3y = 2-2x - 3y/-3 = (2-2x)/-3 y = -2/3+2/3x

OpenStudy (anonymous):

Graph this y = -2/3+2/3x and y = 2x + 2S

OpenStudy (anonymous):

i got -19.25 and y=-13.5

OpenStudy (anonymous):

x=-19.25

OpenStudy (anonymous):

whoops I typed the wrong equation y = -2/3+2/3x and y = 2x + 2

OpenStudy (anonymous):

Intersection should be smaller

OpenStudy (anonymous):

what is it i cant get it

OpenStudy (anonymous):

-2 and -2

OpenStudy (anonymous):

x and y?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so for my final answer it goes x=-2x-2